### v1

# KB-DMGen: Knowledge-Based Global Guidance and Dynamic Pose Masking for Human Image Generation

Abstract
Recent methods using diffusion models have made significant progress in human image generation with various control signals such as pose priors. In portrait generation, both the accuracy of human pose and the overall visual quality are crucial for realistic synthesis. Most existing methods focus on controlling the accuracy of generated poses, but ignore the quality assurance of the entire image. In order to ensure the global image quality and pose accuracy, we propose Knowledge-Based Global Guidance and Dynamic pose Masking for human image Generation (KB-DMGen). The Knowledge Base (KB) is designed not only to enhance pose accuracy but also to leverage image feature information to maintain overall image quality. Dynamic Masking (DM) dynamically adjusts the importance of pose-related regions. Experiments demonstrate the effectiveness of our model, achieving new state-of-the-art results in terms of AP and CAP on the HumanArt dataset. The code will be made publicly available.
Index Terms:
Human Image Generation, knowledge base, dynamic masking
Figure 1: Overview of our method. The visual Knowledge Base (KB) provides global guidance, while the Dynamic pose Mask (DM) enables fine-grained local control. Together, they achieve unified control of both global semantics and local details.
I Introduction
The goal of human image generation (HIG) is to synthesize high-quality images under certain conditions based on a series of prompts (e.g., pose) [1, 2, 3, 4]. Human image generation serves a wide range of real-world applications, including animation [5], game production [6], and other fields.
Previous methods [7, 8, 9] require a source image during training using variational autoencoders (VAEs) [10] or Generative Adversarial Networks (GANs) [11] for dictating the style of the generated images. These methods synthesize target images with specific human features by adjusting the source images, but the training process of these methods is unstable and highly dependent on the distribution of the source images. In contrast, recent advances in controllable text-to-image (T2I) Stable Diffusion (SD) [12] show potential to eliminate the need for source images, enabling greater creative freedom through reliance on text prompts and external conditions [3, 13, 4, 14].These methods often face challenges in accurately matching conditional images with sparse representations such as skeleton pose data [2]. To achieve accurate pose control, various pose guided T2I methods are proposed, such as introducing pose heatmap supervision loss [1], establishing a graph topological structure between the pose priors and latent representation of diffusion models [15], and applying pose masks [2]. These strategies effectively guide the network to focus on pose regions, thereby improving pose fidelity.
While precise pose alignment is essential, high-quality human image generation also demands coherent global structure and semantics. However, these methods [15, 1, 2] emphasize the modeling of pose details while neglecting global structure and semantics, which can easily lead to a decline in the overall quality of the images. To address this issue, we propose Knowledge-Based Global Guidance and Dynamic Masking for Human Image Generation (KB-DMGen). It introduces a visual knowledge base to provide global semantic guidance during image generation, maintaining image quality while improving pose accuracy. Meanwhile, the dynamic masking mechanism adaptively adjusts the weights of pose-related regions, enabling the model to better balance local pose precision. In summary, the contributions of this paper are as follows:
  • Designing a visual knowledge base to improve pose generation accuracy while preserving image quality.
  • Designing dynamic masks to control the precise generation of poses.
Figure 2: The architecture of KB-DMGen. The text encoding is used to retrieve semantic codebook features from a visual Knowledge Base (KB), providing global semantic guidance; meanwhile, pose information generates temporally dynamic masks through the diffusion process, enabling precise control over human pose.
Figure 3: (a) The codebook training. (b) The classifier training .
II Related work
Pose-Guided Human Image Generation. Previous pose-guided approaches take a source image and pose as input to generate photo-realistic human images with preserved appearance, relying on GANs/VAEs for conditional generation [7, 9, 8, 16, 17]. Recently, Bhunia et al. [18] introduce texture diffusion modules and disentangled classifier-free guidance to accurately model appearance-pose relationships and ensure input-output consistency. Zhang et al. [19] propose a novel Dual-task Pose Transformer Network (DPTN), which introduces an auxiliary task (i.e., source-tosource task) and exploits the dual-task correlation to promote the performance of PGPIG. Shen et al. [20] presented a Progressive Conditional Diffusion Model (PCDM), which narrows the gap between character images under target poses and source poses step by step via a three-stage procedure. The above methods [18, 19, 20] remain highly dependent on the distribution of source images, operating under the paradigm of pose transfer that necessitates original image inputs. In contrast, recent work based on pose guided T2I dispenses with the need for source images entirely: HumanSD [1] enhances pose accuracy via heatmap-guided losses and StablePose [2] employs coarse-to-fine masking for precise alignment. GRPose [15] establish a graph topological structure between the pose priors and latent representation of diffusion models to capture the intrinsic associations between different pose parts. However, these methods [1, 2, 15] prioritize pose fidelity and lack global image quality assurance, prompting us to introduce a visual knowledge base for holistic realism and a dynamic mask for pose precision.
Controllable Diffusion Models. Large-scale T2I diffusion models [12, 21, 22, 23, 24] excel in generating diverse high-quality images but lack precision with text-only prompts. Recent studies have enhanced their controllability using conditions such as edges, sketches, and human poses [25, 3, 4, 14, 13, 1, 26, 27], primarily through two approaches: training full T2I models (e.g., Composer [25] constructing diffusion models via decomposition-composition for multi-control, HumanSD [1] fine-tuning SD with pose-specific losses) or developing plug-in adapters for pre-trained models (e.g., T2I-Adapter [4] and GLIGEN [14] integrating lightweight adapters into frozen SD, ControlNet [3] encoding conditions via trainable SD encoder copies, Uni-ControlNet [13] enabling multi-scale conditional injection, and ControlNet++ [26] optimizing cycle consistency). Our approach aligns with the adapter-based paradigm, freezing the pre-trained SD model.
Relation to VQVAE [28]. VQ-VAE [28, 29, 30] introduced the concept of learning discrete latent representations by mapping continuous features into a codebook of quantized vectors. This foundational framework paved the way for subsequent advancements in generating structured, discrete token spaces for various tasks. VQVAE2 [31] uses Vector Quantized Variational AutoEncoder (VQVAE) for large scale image generation. However, the pixel-wise sampling by auto-regressive models is time-consuming. Building on this, Zero-shot T2I [21] explores using a VQVAE-based framework to generate images from text descriptions. Text2Human [32] also uses a VQVAE-style codebook for modeling local clothing attributes to in human image generation. In contrast, we focus on aligning language and vision through a shared discrete token space. Specifically, image features are discretized to form a visual semantic codebook, and a classifier is trained to retrieve relevant tokens based on textual prompts, providing more efficient and interpretable semantic guidance for image generation.
III Method
This section presents the overall framework, followed by detailed training of the codebook and classifier in the knowledge base, and concludes with the dynamic pose masking strategy.
III-A Overview of KB-DMGen
Our goal is to generate high-quality human images conditioned on pose priors. As shown in Fig. 2, our proposed KB-DMGen framework integrates a Diffusion Pipeline, a Knowledge-Based Adapter, and a Dynamic pose Mask Adapter. The language-driven KB Adapter improves semantic alignment and image quality, while the DM Adapter enhances pose accuracy through refined structural guidance.
We adopt Stable Diffusion [12] as our base model, which includes a Variational Autoencoder (VAE) with encoder $E$ and decoder $D$, and a U-Net for denoising and image refinement in the latent space. Given a pose condition $c_{p}\in\mathbb{R}^{H\times W\times C}$ and a text prompt, we use a CLIP text encoder [33] to extract the text embedding $c_{t}$.
The input image is first mapped into a latent representation $z_{0}$ via the VAE encoder. The diffusion process is then guided by the text and pose conditions to iteratively denoise $z_{t}$ using the U-Net, aiming to approximate the data distribution $p(x)$ through maximizing the log-likelihood:
$\mathcal{L}_{d}=\mathbb{E}_{\mathbf{z}_{t},\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\epsilon_{t}-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{c}_{t},\mathbf{c}_{p},t)\right\|^{2}\right]$ (1)
where $z_{t}$ denotes the latent representation at time step $t$, $\mathcal{N}$ denotes the standard normal distribution, $\epsilon_{\theta}$ denotes the noise prediction network and $\epsilon_{t}$ denotes real noise.
The KB-DMGen adapter is a trainable module (KB Frozen) consisting of two components. The Text-query KB extracts visual codebook features from the KB that align with the text prompt, guiding image generation. The DM transforms pose into an attention mask that dynamically adjusts with timesteps. Together, these components guide the diffusion model to generate high-quality images $\hat{x}$.
III-B Knowledge Base
Codebook training The goal of Codebook Training is to learn a discrete representation of image features that can be queried by text. The codebook consists of trainable entries $e={e_{1},e_{2},\dots,e_{K}}$, where $e\in\mathbb{R}^{K\times C}$, with $K$ denoting the number of codebook entries and $C$ the feature dimension.
Given an image, we encode it using CLIP to obtain dense image features $z_{e}\in\mathbb{R}^{N\times C}$, where $N=H\times W$ is the number of spatial tokens and $C$ is the feature dimension. We then compute the squared Euclidean distance between each token in $z_{e}$ and the codebook entries:
$\mathbf{D_{\{i,j\}}}=\|z^{(i)}_{e}\|^{2}+\|e_{j}\|^{2}-2z^{(i)}_{e}e_{j}^{T},$ (2)
where $z_{e}^{(i)}\in\mathbb{R}^{C}$ is the $i$-th image token feature, $e_{j}\in\mathbb{R}^{C}$ is the $j$-th codebook entry, and $\text{sim}_{i,j}$ denotes the squared distance between token $i$ and codebook entry $j$.
This computation results in a similarity matrix $D_{\{i,j\}}\in\mathbb{R}^{N\times K}$, where each row corresponds to an image token and each column to a codebook entry. We then find the nearest codebook entry for each token via:
$a_{i}=\arg\min_{j}\mathbf{D}_{i,j},\quad i=1,\ldots,N,$ (3)
resulting in an index vector $a\in\mathbb{R}^{N}$, where $j\in\{1,2,\cdots,K\}$ and $a_{i}$ represents the codebook index that best matches the $i$th image position element indicates the nearest codebook vector. We then convert the indices $a$ into a one-hot encoding matrix $O\in\mathbb{R}^{N\times K}$, where:
$O_{ij}=\begin{cases}1&\text{if }j=a_{i},\\ 0&\text{otherwise}.\end{cases}$ (4)
Finally, the quantized feature representation $\mathbf{z}_{q}\in\mathbb{R}^{N\times C}$ is obtained by a matrix multiplication between the one-hot matrix and the codebook:
$z_{q}=O\cdot e,$ (5)
where $z_{q}\in\mathbb{R}^{N\times C}$ and $\cdot$ is matrix multiplication. To train the codebook, we use the reconstruction loss:
$\mathcal{L}_{\text{mse}}=\text{MSE}(z_{q},z_{e}).$ (6)
Classifier training The second stage of knowledge base construction trains a classifier to align text features with the learned discrete visual tokens (codebook entries). Given an input text, we extract a global embedding $z_{t}\in\mathbb{R}^{C}$ using a pretrained text encoder (e.g., CLIP). The text embedding $t\in\mathbb{R}^{C}$ is mapped to token logits via a projection:
$y=\text{MLP}(z_{t}),$ (7)
where $y\in\mathbb{R}^{N\times K}$. We supervise the prediction using the discrete indices $a$ obtained from codebook training, optimizing the cross-entropy loss:
$\mathcal{L}_{\text{cls}}=\text{CE}(y,a).$ (8)
This learning process encourages the alignment of textual features with the corresponding visual codebook entries, thus enabling text-conditioned visual token retrieval.
III-C Dynamic Masking
To guide the attention mechanism towards human body regions while suppressing background interference, we introduce a dynamic masking strategy. A binary mask $m\in\{0,1\}^{B\times H\times W}$ is derived from the input pose image, where 1 denotes foreground (human) and 0 indicates background.
Given a diffusion timestep embedding $e\in\mathbb{R}^{B\times C}$, we apply a lightweight MLP to generate a soft gating value:
$g=\text{MLP}(e)\in(0,1).$ (9)
We construct a soft mask $\tilde{m}$ by combining the binary mask with the gating value:
$\tilde{m}=(1+g)\cdot m,$ (10)
[…]
This soft mask is directly applied to the attention logits via element-wise multiplication (Hadamard product, denoted by $\odot$) before the softmax:
$\text{Attn}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\odot\tilde{m}\right)V,$ (11)
where $\odot$ indicates element-wise multiplication between the attention map and the soft mask $\tilde{m}$. It enables dynamic modulation of pose attention weights by adjusting the mask weights over diffusion timesteps.
TABLE I: Results on Human-Art dataset. Methods with * are evaluated on released checkpoints.
Dataset Method Pose Accuracy Image Quality T2I Alignment
[…]
Human-Art SD* [12] 0.24 55.71 2.30 11.53 3.36 33.33
T2I-Adapter [4] 27.22 65.65 1.75 11.92 2.73 33.27
ControlNet [3] 39.52 69.19 1.54 11.01 2.23 32.65
Uni-ControlNet [13] 41.94 69.32 1.48 14.63 2.30 32.51
GLIGEN [14] 18.24 69.15 1.46 32.52
HumanSD [1] 44.57 69.68 1.37 10.03 2.70 32.24
GRPose [15] 49.50 70.84 1.43 13.76 2.53 32.31
Stable-Pose [2] 48.88 70.83 1.50 11.12 2.35 32.60
KB-DMGen 51.71 71.40 1.53 10.29 2.45 32.45
Figure 4: The visualization results of KB-DMGen.
TABLE II: Ablation experiments on Human-Art dataset
[…]
+KB 50.73 71.04 1.58 11.28 2.52 32.47
+KB+DM 51.40 71.17 1.54 10.56 2.54 32.41
+KB+DM+ D&C 51.71 71.40 1.53 10.29 2.45 32.45
IV Experiments
IV-A Experimental Settings
Datasets We evaluated our model on the Human-Art [34] dataset. The Human-Art dataset comprises 50,000 high-quality images from 5 real-world and 15 virtual scenarios, featuring human bounding boxes, key points and textual descriptions.
Implementation Details Similar to previous work [3, 4, 2], we fine-tune our model on SD with version 1.5. We utilize Adam [35] optimizer with a learning rate of $1\times 10^{−5}$. We also follow [3] to randomly replace text prompts as empty strings at a probability of 0.5, which aims to strengthen the control of the pose input. During inference, no text prompts are removed and a DDIM sampler [36] with 50 time steps is utilized to generate images. We train our models with 10 epochs.
For the Knowledge Base, we fine-tune on the OpenCLIP [37, 38] implementation using the pretrained CLIP model (ViT-L/14) with frozen encoders. Training is performed in two stage—codebook optimization and classifier fine-tuning—both using the same strategy over 30 epochs with AdamW (initial learning rate 0.001) and a cosine decay schedule. The training and validation splits as same as Human-Art [34].
Metrics. To evaluate pose accuracy, we use mean Average Precision (AP), Pose Cosine Similarity-based AP (CAP), and People Counting Error (PCE) [39], computed using HigherHRNet [40] to compare ground-truth poses with those extracted from generated images. For image quality assessment, we employ Fréchet Inception Distance (FID) [41] and Kernel Inception Distance (KID) [42] to measure diversity and fidelity. Text-image alignment is evaluated using the CLIP score [33].
IV-B Comparison with SOTA Methods
Quantitative results comparing our method with other state-of-the-art (SOTA) approaches are shown in Table I.
The experimental results indicate that KB-DMGen achieved the highest AP and CLIP-score. On the Human-Art dataset, our model attained AP and CAP of 51.71% and 71.40% respectively, bringing an improvement of 2.21% in AP, compared with GRPose [15] and an improvement of 0.56% in CAP.Notably, our method significantly improves pose accuracy while ensuring the global quality (as measured by FID and KID). KID is multiplied by 100 for Human-Art.
We further compared the visualized qualitative results with other methods. As shown in Fig. 4, our method is superior in terms of image pose accuracy and the matching degree between images and text.
IV-C Ablation Studies
Ablation experiments on the Human-Art dataset to validate the effectiveness of each module, as shown in Table II.
KB. When introducing the visual knowledge base, our method outperforms the recent SOTA method GRPose [15] with a 1.23 AP and 0.2 CAP improvement, while achieving better image quality with a 2.48 drop in FID and a slight 0.01 decrease in KID. This demonstrates that the KB enhances pose accuracy without compromising and even improving image quality.
DM. Building on the knowledge base, we further introduce Dynamic Masking (DM), which brings a 0.67 improvement in pose accuracy and a 0.72 reduction in FID, indicating enhanced image quality.
Hierarchical Language. To enhance the semantic parsing capability of the classifier, we decompose the language input into individual components, then query the codebook from the knowledge base (KB) for each component separately. The queried results are fused and injected into the stable diffusion (SD) model. We refer to this process as ”Decomposition and Combination (D&C)”. The results show that adding D&C based on KB and DM further improves AP and SAP, with other metrics also optimized to varying degrees. This demonstrates the effectiveness of D&C.
V Conclusion
We present an image generation approach that integrates a visual knowledge base with dynamic masking. Experimental results demonstrate the effectiveness of the proposed method.